# Clairaut anti-invariant Riemannian map from/to Kähler manifolds admitting Ricci solitons

### v1

@@ 2 Preliminaries
[12] Let $F:(N,g_{N})\rightarrow(M,g_{M})$ be a proper Riemannian map from a Riemannian manifold to an almost Hermitian manifold with almost complex structure $J.$ We say that $F$ is an anti-invariant Riemannian map at $q\in N$ if $J(rangeF_{*}q)\subset(rangeF_{*q})^{\perp}.$
If $F$ is an anti-invariant Riemannian map for every $q\in N,$ then $F$ is called an anti-invariant Riemannian map. In this case, we denote the orthogonal subbundle to $J(rangeF_{*})$ in $(rangeF_{*})^{\perp}$ by $\nu$, i.e., $(rangeF_{*})^{\perp}=J(rangeF_{*})\oplus\nu.$
For any $E\in\Gamma(rangeF_{*})^{\perp},$ we have [19]
$JE=PE+QE,$ (2.8)
where $PE\in\Gamma(rangeF_{*})$ and $QE\in\nu$. Note that if $\nu=0$ then $F$ is called a Lagrangian Riemannian map.
Theorem 2.4.
[7] Let $F:(M,g_{M})\rightarrow(N,g_{N})$ be a Clairaut Riemannian map with $\tilde{s}=e^{g}$ between Riemannian manifolds such that $kerF_{*}$ is minimal. Then $F$ is harmonic if and only if g is a constant function on N.
@@
In this section, we study Clairaut anti-invariant Riemannian maps from Kähler manifolds admitting Ricci solitons to Riemannian manifolds. Throughout this section, we are taking $dim(kerF_{*})>1.$ Below we provide a revised version of proposition (3.9) [22].
@@
$\begin{split}Ric(X,Y)&=Ric^{(KerF_{*})}(BX,BY)-\Big{(}r|\nabla f|^{2}+div(\nabla f)\Big{)}g_{M}(BX,BY)\\ &+\sum_{i=1}^{r+s}g_{M}(A_{X_{i}}BX,A_{X_{i}}BY)-rg_{M}(\nabla_{CX}\nabla f,CY)\\ &-rCX(f)CY(f)+\sum_{j=1}^{r}g_{M}(A_{CX}u_{j},A_{CY}u_{j})\\ &+div\Big{(}A(CX,CY)\Big{)}+Ric^{(rangeF_{*})}(F_{*}CX,F_{*}CY)\\ &-\sum_{i=1}^{r+s}g_{N}\Big{(}(\nabla F_{*})(X_{i},CY),(\nabla F_{*})(CX,X_{i})\Big{)}+g_{N}\Big{(}(\nabla F_{*})(CX,CY),\tau^{(kerF_{*})^{\perp}}\Big{)}\\ &-(r+1)g_{M}(\nabla_{BX}\nabla f,CY)-\sum_{i=1}^{r+s}g_{M}\Big{(}(\nabla_{X_{i}}A)(CY,X_{i}),BX\Big{)}\\ &-(r+1)g_{M}(\nabla_{BY}\nabla f,CX)-\sum_{i=1}^{r+s}g_{M}\Big{(}(\nabla_{X_{i}}A)(CX,X_{i}),BY\Big{)}\end{split}$ (3.3)
@@
$\begin{split}Ric(U,V)=&\sum_{j=1}^{r}\Big{\{}-g_{M}\Big{(}\nabla_{JU}T)(u_{j},u_{j}),JV\Big{)}-g_{M}\Big{(}(\nabla_{u_{j}}A)(JU,JV),u_{j}\Big{)}\\ &+g_{M}(T_{u_{j}}JU,T_{u_{j}}JV)-g_{M}(A_{JU}u_{j},A_{JV}u_{j})\Big{\}}\\ &+Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)+g_{N}\Big{(}(\nabla F_{*})(JU,JV),\tau^{(kerF_{*})^{\perp}}\Big{)}\\ &-\sum_{i=1}^{r+s}g_{N}\Big{(}(\nabla F_{*})(X_{i},JV),(\nabla F_{*})(JU,X_{i})\Big{)}.\end{split}$ (3.4)
Since $F$ is a Clairaut anti-invariant Riemannian map, using clairaut condition $T_{U}V=-g(U,V)\nabla f,$ Lemma 3.7 [22] and Lemma 3.8 [22], we get
$-\sum_{j=1}^{r}g_{M}\Big{(}\nabla_{JU}T)(u_{j},u_{j}),JV\Big{)}=rg_{M}(\nabla_{JU}\nabla f,JV),$
@@
$Ric(U,V)=Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)+rg_{M}(\nabla_{JU}\nabla f,JV)-div\Big{(}A(JU,JV)\Big{)},$
@@
$\begin{split}Ric(U,X)=&\sum_{j=1}^{r}\Big{\{}g_{M}\Big{(}(\nabla_{BX}T)(u_{j},u_{j}),JU\Big{)}-g_{M}\Big{(}(\nabla_{u_{j}}T)(BX,u_{j}),JU\Big{)}\\ &+g_{M}\Big{(}(\nabla_{JU}T)(u_{j},u_{j}),CX\Big{)}+g_{M}\Big{(}(\nabla_{u_{j}}A)(JU,CX),u_{j}\Big{)}\\ &-g_{M}(T_{u_{j}}JU,T_{u_{j}}CX)+g_{M}(A_{JU}u_{j},A_{CX}u_{j})\Big{\}}+Ric^{(rangeF_{*})}(F_{*}JU,F_{*}CX)\\ &+g_{N}\Big{(}(\nabla F_{*})(JU,CX),\tau^{(kerF_{*})^{\perp}}\Big{)}+\sum_{i=1}^{r+s}\Big{\{}g_{M}\Big{(}(\nabla_{X_{i}}A)(X_{i},JU),BX\Big{)}\\ &+2g_{M}(A_{X_{i}}JU,T_{BX}X_{i})-g_{N}\Big{(}(\nabla F_{*})(X_{i},CX),(\nabla F_{*})(JU,X_{i}))\Big{\}}.\end{split}$ (3.5)
Since $F$ is a Clairaut anti-invariant Riemanian map, using clairaut condition $T_{U}V=-g(U,V)\nabla f,$ lemma 3.7 [22] and lemma 3.8 [22], we get
@@
$Ric(X,Y)=Ric(JX,JY)~{}~{}\forall~{}~{}X,Y\in\Gamma(TM).$ (3.6)
@@
using lemma 52 ([16]), we get
$\begin{split}Ric(BX,BY)&=Ric^{(BX,BY)}-rg_{M}(H,T_{BX}BY)+g_{M}(A_{X_{i}}BX,A_{X_{i}}BY)\\ &+\sum_{i=1}^{r+s}g_{M}\Big{(}(\nabla_{X_{i}}T)_{BX}BY,X_{i}\Big{)}.\end{split}$
@@
$\begin{split}Ric(BX,BY)&=Ric^{(BX,BY)}-r|\nabla f|^{2}g_{M}(BX,BY)+\sum_{i=1}^{r+s}g_{M}(A_{X_{i}}BX,A_{X_{i}}BY)\\ &-g_{M}(BX,BY)div(\nabla f).\end{split}$ (3.8)
@@
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a Clairaut Langragian Riemannian map from a Kähler manifold to a Riemannian manifold with $\tilde{r}=e^{f}.$ Then, Ricci tensors on $(M,g_{M})$ are given by
$\begin{split}Ric(U,V)=rg_{M}(\nabla_{JU}\nabla f,JV)+Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV),\end{split}$ (3.12)
@@
$\begin{split}Ric(X,Y)=Ric^{(KerF_{*})}(BX,BY)-\Big{(}r|\nabla f|^{2}+div(\nabla f)\Big{)}g_{M}(BX,BY).\end{split}$ (3.14)
@@
In case of Langragian Riemannian map, $\mu=0$ or $CX=0=CY.$
@@
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a totally geodesic Clairaut anti-invariant Riemannian map from a Kähler manifold to a Riemannian manifold with with $\tilde{r}=e^{f}$ such that $(kerF_{*})^{\perp}$ integrable. Then
$\begin{split}Ric(X,Y)&=Ric^{(kerF_{*})}(BX,BY)-\Big{(}r|\nabla f|^{2}+div(\nabla f)\Big{)}g_{M}(BX,BY)\\ &-rg_{M}(\nabla_{CX}\nabla f,CY)+Ric^{(rangeF_{*})}(F_{*}CX,F_{*}CY)-rCX(f)CY(f),\end{split}$ (3.15)
@@
Since $F$ is a totally geodesic Clairaut anti-invariant Riemannian map, the second fundamental form of the map $\nabla F_{*}$ vanishes. Since $(kerF_{*})^{\perp}$ is integrable, we have $A_{X}=0~{}~{}\forall X\in\Gamma(kerF_{*})^{\perp}$. Therefore the equation (3.3), reduces to (3.15). ∎
@@
Since $(M,g_{M},J)$ is a Kähler manifold admitting a Ricci soliton with potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$
@@
If $\eta=W\in\Gamma(kerF_{*})$ or in $\Gamma(\mu)$, then above equation gives
@@
If $\eta\in\Gamma(KerF_{*})^{\perp}$ but $\notin\Gamma(\mu)$, then from (3.18), we have
@@
Using Theorem 2.3 and Theorem 3.5 ([22]) , we get
@@
Let $F$ be a Clairaut Langrangian Riemannian map from a Kähler manifold $(M,g_{M},J)$ admitting Ricci soliton with potential vector field $\eta\in(TM)$ and scalar $\lambda$ to a Riemannian manifold $(N,g_{N})$ with $r=e^{f}$. Then $rangeF_{*}$ is an $\alpha$- Ricci soliton.
@@
Using Theorem 3.1 and Lemma 56 ([14]), we get the required result. ∎
@@
Let $F$ be a Clairaut anti-invariant Riemannian map from a Kähler manifold $(M,g_{M},J)$ admitting Ricci soliton with potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$ to an Einstein manifold $(N,g_{N})$. Then $\nabla f$ is a conformal vector field on $J(kerF_{*}).$
@@
Since $rangeF_{*}$ is Einstein, $Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)=\lambda g_{N}(F_{*}JU,F_{*}JV).$
@@
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a Clairaut Langrangian Riemannian map from a Kähler manifold admitting Ricci soliton with potential vector fields $\eta\in\Gamma(TM)$ and scalar $\lambda$ to a Riemannian manifold. Then $s^{(rangeF_{*})}=-rdiv(\nabla f)-\lambda r.$
@@
If $\eta$ is either horizontal or vertical(as we have seen in theorem 6), we get
@@
$\begin{split}&rg_{M}(\nabla_{JU}\nabla f,JV)+Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)\\ &+\lambda g_{M}(U,V)=0.\end{split}$
@@
$rdiv(\nabla f)+s^{(rangeF_{*})}+\lambda r=0,$
which implies
$s^{(rangeF_{*})}=-rdiv(\nabla f)-\lambda r.$ (3.20)
@@
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a Clairaut Langragian Riemannian map from a Kähler manifold admitting Ricci soliton with vertical potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$ to a Riemannian manifold. Then $kerF_{*}$ is an Einstein.
@@
$\begin{split}&\dfrac{1}{2}(L_{\eta}g_{M})(X,Y)+Ric^{(KerF_{*})}(BX,BY)-\Big{(}r|\nabla f|^{2}+div(\nabla f)\Big{)}g_{M}(BX,BY)\\ &+\lambda g_{M}(X,Y)=0,\end{split}$
@@
$\begin{split}&\dfrac{1}{2}\Big{(}g_{M}(\nabla_{X}{\eta},Y)+g_{M}(\nabla_{Y}{\eta},X)\Big{)}-\Big{(}r|\nabla f|^{2}+div(\nabla f)\Big{)}g_{M}(BX,BY)\\ &+Ric^{(KerF_{*})}(BX,BY)+\lambda g_{M}(JX,JY)=0.\end{split}$
Since $\eta$ is vertical, above equation gives
$Ric^{(KerF_{*})}(BX,BY)+\nu^{\prime}g_{M}(BX,BY)=0,$ (3.21)
where $\nu^{\prime}=-r|\nabla f|^{2}-div(\nabla f)+\lambda$ is a function on $M.$ This implies that $(kerF_{*})$ is an Einstein.
@@
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a Clairaut Langragian Riemannian map from a Kähler manifold admitting Ricci soliton with vertical potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$ to a Riemannian manifold. Then scalar curvature $s^{kerF_{*}}=-r\nu^{\prime}.$
@@
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a totally geodesic Clairaut anti-invariant Riemannian map from a Kähler manifold admitting Ricci soliton with vertical potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$ to a Riemannian manifold such that $(kerF_{*})^{\perp}$ is integrable and $f$ is constant on $\Gamma(\mu).$ Then $kerF_{*}$ is an Einstein, provided $F_{*}(\mu)$ is an Einstein.
@@
Since $(M,g_{M},J)$ is a Kähler manifold admitting a Ricci soliton with vertical potential vector field $\eta\in\Gamma(TM)$. Then, we have
@@
$\begin{split}&\dfrac{1}{2}\Big{(}g_{M}(\nabla_{X}\eta,Y)+g_{M}(\nabla_{Y}\eta,X)\Big{)}+Ric^{(kerF_{*})}(BX,BY)-rg_{M}(\nabla_{CX}\nabla f,CY)\\ &-(r|\nabla f|^{2}+div(\nabla f)-\lambda)g_{M}(BX,BY)+Ric^{(rangeF_{*})}(F_{*}CX,F_{*}CY)\\ &-rCX(f)CY(f)+\lambda g_{N}(F_{*}CX,F_{*}CY)=0.\end{split}$
Since $(kerF_{*})^{\perp}$ is integrable and $f$ is constant on $\Gamma(\mu),$ from above equation, we get
$\begin{split}&Ric^{(kerF_{*})}(BX,BY)+\lambda g_{M}(BX,BY)+Ric^{(rangeF_{*})}(F_{*}CX,F_{*}CY)+\lambda g_{N}(F_{*}CX,F_{*}CY)=0.\end{split}$
Since $F_{*}(\mu)$ is an Einstein, then above equation reduces to
$\begin{split}Ric^{(kerF_{*})}(BX,BY)+\lambda g_{M}(BX,BY)=0.\end{split}$
Therefore from above equation, we can say that $F_{*}(\mu)$ is an Einstein. ∎

### v2

@@ 2.2 Some results on Clairaut anti-invariant Riemannian maps to Kähler manifolds
[11] Let $F:(M,g_{M})\rightarrow(N,g_{N})$ be a Riemannian map between Riemannian manifolds such that $(rangeF_{*})^{\perp}$ is totally geodesic and $rangeF_{*}$ is connected, let $\beta$ and $\sigma=Fo\beta$ be geodesics on $M$ and $N,$ respectively. Then $F$ is a Clairaut Riemannian map with $\tilde{s}=e^{g}$ if and only if any one of the following conditions holds:
  • (i)
    $S_{D}F_{*}X=-D(g)F_{*}X$, where $F_{*}X\in\Gamma(rangeF_{*})$ and $D\in\Gamma(rangeF_{*})^{\perp}$ are components of $\dot{\sigma}(t)$.
  • (ii)
    $F$ is an umbilical map, and has $H^{\prime}=-\nabla^{N}g$ , where $g$ is a smooth function on $N$ and $H^{\prime}$ is the mean curvature vector field of $rangeF_{*}$.
@@
[20] Let $F:(M,g_{M})\rightarrow(N,g_{N},J^{\prime})$ be a proper Riemannian map from a Riemannian manifold to an almost Hermitian manifold with almost complex structure $J^{\prime}.$ We say that $F$ is an anti-invariant Riemannian map at $q\in N$ if $J^{\prime}(rangeF_{*}q)\subset(rangeF_{*q})^{\perp}.$
If $F$ is an anti-invariant Riemannian map for every $q\in N,$ then $F$ is called an anti-invariant Riemannian map. In this case, we denote the orthogonal subbundle to $J^{\prime}(rangeF_{*})$ in $(rangeF_{*})^{\perp}$ by $\nu$, i.e., $(rangeF_{*})^{\perp}=J^{\prime}(rangeF_{*})\oplus\nu.$
For any $D\in\Gamma(rangeF_{*})^{\perp},$ we have [27]
$J^{\prime}D=PD+QD,$ (2.8)
where $PD\in\Gamma(rangeF_{*})$ and $QD\in\Gamma(\nu)$. Note that if $\nu=0$ then $F$ is called a Lagrangian Riemannian map.
Now, we recall the definition of Clairaut anti-invariant Riemannian map for target manifold.
Definition 2.6.
[11] An anti-invariant Riemannian map from a Riemannian manifold to a Kähler manifold is called Clairaut anti-invariant Riemannian map if it satisfies the definition of Clairaut Riemannian manifold for target manifold.
@@
In this section, we study Clairaut anti-invariant Riemannian maps from Kähler manifolds admitting Ricci solitons to Riemannian manifolds. Throughout this section, we are taking $dim(kerF_{*})>1.$ Now, we improve the Proposition (3.9) [30] to explore the study in our case.
@@
$\begin{split}Ric(X,Y)&=Ric^{(KerF_{*})}(BX,BY)-\Big{(}r||\nabla f||^{2}+div(\nabla f)\Big{)}g_{M}(BX,BY)\\ &+\sum_{i=1}^{r+s}g_{M}(A_{X_{i}}BX,A_{X_{i}}BY)-rg_{M}(\nabla_{CX}\nabla f,CY)\\ &-rCX(f)CY(f)+\sum_{j=1}^{r}g_{M}(A_{CX}u_{j},A_{CY}u_{j})\\ &+div\Big{(}A(CX,CY)\Big{)}+Ric^{(rangeF_{*})}(F_{*}CX,F_{*}CY)\\ &-\sum_{i=1}^{r+s}g_{N}\Big{(}(\nabla F_{*})(X_{i},CY),(\nabla F_{*})(CX,X_{i})\Big{)}+g_{N}\Big{(}(\nabla F_{*})(CX,CY),\tau^{(kerF_{*})^{\perp}}\Big{)}\\ &-(r+1)g_{M}(\nabla_{BX}\nabla f,CY)-\sum_{i=1}^{r+s}g_{M}\Big{(}(\nabla_{X_{i}}A)(CY,X_{i}),BX\Big{)}\\ &-(r+1)g_{M}(\nabla_{BY}\nabla f,CX)-\sum_{i=1}^{r+s}g_{M}\Big{(}(\nabla_{X_{i}}A)(CX,X_{i}),BY\Big{)}\end{split}$ (3.3)
@@
$\begin{split}Ric(U,V)=&\sum_{j=1}^{r}\Big{\{}-g_{M}\Big{(}(\nabla_{JU}T)(u_{j},u_{j}),JV\Big{)}-g_{M}\Big{(}(\nabla_{u_{j}}A)(JU,JV),u_{j}\Big{)}\\ &+g_{M}(T_{u_{j}}JU,T_{u_{j}}JV)-g_{M}(A_{JU}u_{j},A_{JV}u_{j})\Big{\}}\\ &+Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)+g_{N}\Big{(}(\nabla F_{*})(JU,JV),\tau^{(kerF_{*})^{\perp}}\Big{)}\\ &-\sum_{i=1}^{r+s}g_{N}\Big{(}(\nabla F_{*})(X_{i},JV),(\nabla F_{*})(JU,X_{i})\Big{)}.\end{split}$ (3.4)
Since $F$ is a Clairaut anti-invariant Riemannian map, using Clairaut condition $T_{U}V=-g(U,V)\nabla f,$ Lemmas 3.7 and 3.8 of [30], we get
$-\sum_{j=1}^{r}g_{M}\Big{(}(\nabla_{JU}T)(u_{j},u_{j}),JV\Big{)}=rg_{M}(\nabla_{JU}\nabla f,JV),$
@@
$Ric(U,V)=Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)+rg_{M}(\nabla_{JU}\nabla f,JV)-div\Big{(}A(JU,JV)\Big{)}{\color[rgb]{1,0,0},}$
@@
$\begin{split}Ric(U,X)=&\sum_{j=1}^{r}\Big{\{}g_{M}\Big{(}(\nabla_{BX}T)(u_{j},u_{j}),JU\Big{)}-g_{M}\Big{(}(\nabla_{u_{j}}T)(BX,u_{j}),JU\Big{)}\\ &+g_{M}\Big{(}(\nabla_{JU}T)(u_{j},u_{j}),CX\Big{)}+g_{M}\Big{(}(\nabla_{u_{j}}A)(JU,CX),u_{j}\Big{)}\\ &-g_{M}(T_{u_{j}}JU,T_{u_{j}}CX)+g_{M}(A_{JU}u_{j},A_{CX}u_{j})\Big{\}}+Ric^{(rangeF_{*})}(F_{*}JU,F_{*}CX)\\ &+g_{N}\Big{(}(\nabla F_{*})(JU,CX),\tau^{(kerF_{*})^{\perp}}\Big{)}+\sum_{i=1}^{r+s}\Big{\{}g_{M}\Big{(}(\nabla_{X_{i}}A)(X_{i},JU),BX\Big{)}\\ &+2g_{M}(A_{X_{i}}JU,T_{BX}X_{i})-g_{N}\Big{(}(\nabla F_{*})(X_{i},CX),(\nabla F_{*})(JU,X_{i})\Big{\}}.\end{split}$ (3.5)
Since $F$ is a Clairaut anti-invariant Riemanian map, using Clairaut condition $T_{U}V=-g(U,V)\nabla f,$ Lemmas 3.7 and 3.8 of [30], we get
@@
$Ric(X^{\prime},Y^{\prime})=Ric(JX^{\prime},JY^{\prime})~{}~{}\forall~{}~{}X^{\prime},Y^{\prime}\in\Gamma(TM).$ (3.6)
@@
using Lemma 52 of [24], we get
$\begin{split}Ric(BX,BY)&=Ric^{kerF_{*}}(BX,BY)-rg_{M}(H,T_{BX}BY)+g_{M}(A_{X_{i}}BX,A_{X_{i}}BY)\\ &+\sum_{i=1}^{r+s}g_{M}\Big{(}(\nabla_{X_{i}}T)_{BX}BY,X_{i}\Big{)}.\end{split}$
@@
$\begin{split}Ric(BX,BY)&=Ric^{kerF_{*}}(BX,BY)-r||\nabla f||^{2}g_{M}(BX,BY)+\sum_{i=1}^{r+s}g_{M}(A_{X_{i}}BX,A_{X_{i}}BY)\\ &-g_{M}(BX,BY)div(\nabla f).\end{split}$ (3.8)
@@
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a Clairaut Lagrangian Riemannian map from a Kähler manifold to a Riemannian manifold with $\tilde{r}=e^{f}.$ Then, Ricci tensors on $(M,g_{M},J)$ are given by
$\begin{split}Ric(U,V)=Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV),\end{split}$ (3.12)
@@
$\begin{split}Ric(X,Y)=Ric^{(KerF_{*})}(BX,BY).\end{split}$ (3.14)
@@
In case of Lagrangian Riemannian map, $\mu=0$ or $CX=0=CY$ and from Lemma 56 of [22], we have $A_{X}=0~{}~{}\forall X\in\Gamma(kerF_{*})^{\perp}.$ From Theorem 3.5 of [30], we get $f$ is constant on $(kerF_{*})^{\perp}.$
@@
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a totally geodesic Clairaut anti-invariant Riemannian map from a Kähler manifold to a Riemannian manifold with with $\tilde{r}=e^{f}$ such that $(kerF_{*})^{\perp}$ is totally geodesic. Then
$\begin{split}Ric(X,Y)&=Ric^{(kerF_{*})}(BX,BY)-\Big{(}r||\nabla f||^{2}+div(\nabla f)\Big{)}g_{M}(BX,BY)\\ &-rg_{M}(\nabla_{CX}\nabla f,CY)+Ric^{(rangeF_{*})}(F_{*}CX,F_{*}CY)-rCX(f)CY(f),\end{split}$ (3.15)
@@
Since $F$ is a totally geodesic Clairaut anti-invariant Riemannian map, the second fundamental form of the map $\nabla F_{*}$ vanishes. Since $(kerF_{*})^{\perp}$ is totally geodesic, we have $A_{X}=0~{}~{}\forall X\in\Gamma(kerF_{*})^{\perp}$. Therefore the (3.3), reduces to (3.15). ∎
@@
Since $(M,g_{M},J)$ is a Kähler manifold admitting a Ricci soliton with potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda,$ we have
@@
Case (i): When $\eta\in\Gamma(kerF_{*})$ or $\eta\in\Gamma(\mu).$
In this case (3.18)
@@
Case (ii): When $\eta\in\Gamma(KerF_{*})^{\perp}$ but $\notin\Gamma(\mu).$
In this case from (3.18), we get
@@
Using Theorem 2.1 and Theorem 3.5 of [30] , we get
@@
Therefore, in either case $F_{*}(JkerF_{*})$ is an $\alpha$-Ricci soliton. ∎
@@
Let $F$ be a Clairaut Lagrangian Riemannian map from a Kähler manifold $(M,g_{M},J)$ admitting Ricci soliton with potential vector field $\eta\in(TM)$ and scalar $\lambda$ to a Riemannian manifold $(N,g_{N})$ with $\tilde{r}=e^{f}$. Then $rangeF_{*}$ is an Einstein.
@@
Using Theorem 3.1, Lemma 56 of [22] and Theorem 3.5 of ([30]), we get the required result. ∎
@@
Let $F$ be a Clairaut anti-invariant Riemannian map from a Kähler manifold $(M,g_{M},J)$ admitting Ricci soliton with potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$ to an Einstein manifold $(N,g_{N})$ such that $(kerF_{*})^{\perp}$ is totally geodesic. Then $\nabla f$ is a conformal vector field on $J(kerF_{*}).$
@@
Since $rangeF_{*}$ is an Einstein, $Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)=\lambda g_{N}(F_{*}JU,F_{*}JV).$
@@
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a Clairaut Lagrangian Riemannian map from a Kähler manifold admitting Ricci soliton with potential vector fields $\eta\in\Gamma(TM)$ and scalar $\lambda$ to a Riemannian manifold. Then $s^{(rangeF_{*})}=-\lambda r.$
@@
Using Lemma 56 of [24] and Theorem 3.1, we get $(L_{\eta}g_{M})(U,V)=0.$ Then above equation reduces to
@@
$\begin{split}Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)+\lambda g_{M}(U,V)=0.\end{split}$
@@
$s^{(rangeF_{*})}+\lambda r=0,$
@@
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a Clairaut Lagrangian Riemannian map from a Kähler manifold admitting Ricci soliton with vertical potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$ to a Riemannian manifold. Then $kerF_{*}$ is an Einstein.
@@
$\begin{split}\dfrac{1}{2}(L_{\eta}g_{M})(X,Y)+Ric^{(KerF_{*})}(BX,BY)+\lambda g_{M}(X,Y)=0,\end{split}$
@@
$\begin{split}\dfrac{1}{2}\Big{(}g_{M}(\nabla_{X}{\eta},Y)+g_{M}(\nabla_{Y}{\eta},X)\Big{)}+Ric^{(KerF_{*})}(BX,BY)+\lambda g_{M}(JX,JY)=0.\end{split}$
Since $\eta$ is vertical potential vector field, above equation gives
$Ric^{(KerF_{*})}(BX,BY)+\lambda g_{M}(BX,BY)=0.$ (3.20)
This implies that $kerF_{*}$ is an Einstein.
@@
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a Clairaut Lagrangian Riemannian map from a Kähler manifold admitting Ricci soliton with vertical potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$ to a Riemannian manifold. Then scalar curvature of $kerF_{*}$ is $s^{kerF_{*}}=-r\lambda.$
@@
Next, we give a theorem which shows the relation between Ricci curvature and Riemannian metric.
@@
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a totally geodesic Clairaut anti-invariant Riemannian map from a Kähler manifold admitting Ricci soliton with vertical potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$ to a Riemannian manifold such that $(kerF_{*})^{\perp}$ is totally geodesic and $f$ is constant on $\Gamma(\mu).$ Then
$Ric^{(rangeF_{*})}(F_{*}CX,F_{*}CY)+\lambda g_{N}(F_{*}CX,F_{*}CY)=0$
satisfies provided $kerF_{*}$ is an Einstein.
@@
$\begin{split}&\dfrac{1}{2}\Big{(}g_{M}(\nabla_{X}\eta,Y)+g_{M}(\nabla_{Y}\eta,X)\Big{)}+Ric^{(kerF_{*})}(BX,BY)-rg_{M}(\nabla_{CX}\nabla f,CY)\\ &-(r||\nabla f||^{2}+div(\nabla f)-\lambda)g_{M}(BX,BY)+Ric^{(rangeF_{*})}(F_{*}CX,F_{*}CY)\\ &-rCX(f)CY(f)+\lambda g_{N}(F_{*}CX,F_{*}CY)=0.\end{split}$
Since $(kerF_{*})^{\perp}$ is totally geodesic and $f$ is constant on $\Gamma(\mu),$ from above equation, we get
$\begin{split}&Ric^{(kerF_{*})}(BX,BY)+\lambda g_{M}(BX,BY)+Ric^{(rangeF_{*})}(F_{*}CX,F_{*}CY)\\ &+\lambda g_{N}(F_{*}CX,F_{*}CY)=0.\end{split}$
Since $kerF_{*}$ is an Einstein, then above equation reduces to
$\begin{split}Ric^{(rangeF_{*})}(F_{*}CX,F_{*}CY)+\lambda g_{N}(F_{*}CX,F_{*}CY)=0.\end{split}$
This completes the proof. ∎
Next, we give a theorem which shows the relation between Ricci curvature and Lie derivative.